# WhyFlow: Interrogative Debugger for Sensemaking Taint Analysis

### v2

@@ 3. Speculative Analysis for Inquiry-based Debugging
WhyFlow enables inquiry-based debugging for sensemaking to debug spurious flows or missing flows. As presented in Figure 7, we execute both the CodeQL taint query as well as a less restricted version of this query (Sections 3.1), convert the taint analysis results into Soufflé facts and store them in a database (Section 3.2). Subsequently, WhyFlow queries the database to provide answers to the questions posed using WhyFlow, according to the user’s choice (Section 3.3). The answers are presented as an interactive graph visualization. In this section, we present the components of WhyFlow, and how each of these components help the overall workflow. Before we elaborate on our approach, we state several assumptions in our design.
Assumption 1. We presume that a user of WhyFlow is familiar with the subject program that she is inspecting, meaning that they may recognize that some flows are spurious, unexpected flows and some expected flows are missing, serving a starting point for end-user debugging. This way, running an interrogative debugger can help narrow down the third-party library’s model after selecting a relevant template query for ‘why’ or ‘why-not’ questions.
Assumption 2. We leverage CodeQL’s taint analysis as is obtain taint analysis results, and then we utilize Soufflé to run the interrogative debugging queries on top of CodeQL results. Hence, it is important to emphasize that CodeQL and Soufflé are being used together, and there is no modification to the underlying CodeQL’s analysis algorithm. These two tools are not meant to compete.
Assumption 3. We assume that CodeQL results are created with the information flow models of third party libraries, which is typical for modern taint analysis. Since these models are plug-in abstractions and third party libraries that may not have source code, we assume that inaccuracies in these plug-in models should be debugged, when a user suspects a missing flow or a spurious flow.
Assumption 4. The goal of WhyFlow is not to frame ‘why’ and ‘why-not’ question asking as a classification problem of false positive or false negative. In fact, WhyFlow never removes or adds any taint warnings reported by CodeQL. Rather, WhyFlow helps users with end-user debugging—how their configuration choices such as source/sink/sanitizer definitions and third-party models may lead to spurious or missing paths. Given that the target users need to maintain and update third-party models, WhyFlow aims to simplify this process.
@@ 3.1. CodeQL
CodeQL is an open-source static analysis framework, developed and maintained by GitHub. Users can use a specialized Domain-Specific Language (DSL) to detect potential security vulnerabilities, code quality issues, or other custom patterns (codeql_docs, ). While CodeQL offers code scanning capabilities out of the box, its extendability via user-defined queries and models is particularly noteworthy.
A key factor in CodeQL’s precision and comprehensiveness lies in its third party library modeling. Each library API (including third-party dependencies) can be modeled to reflect how data flows through its methods. Recognizing that manual creation of such models is time-consuming and error-prone, the CodeQL team has explored automated and machine-learning-based techniques to infer library behaviors (nachtigall2022large, ). Nonetheless, incomplete or incorrect modeling remains a practical challenge, especially in large ecosystems where libraries evolve frequently without thorough documentation of their input-output contracts.
Rather than relying on specific third-party library models and source and sink definitions, a general query may relax these configuration assumptions, revealing the “maximal” set of reachable flows from all possible sinks to all possible sources, ignoring third-party library models. Such an approach can help users uncover paths that would otherwise be missed due to incomplete or inaccurate library modeling. Moreover, comparing results from both model-aware and “maximal” queries can expose discrepancies—hints that the current model assumptions may need refinement. Our work builds on these insights by introducing an interactive, interrogative debugger, where users can easily toggle assumptions about sanitizers or third-party methods with pre-defined template queries.
Figure 8. CodeQL’s tree-based view shows each warning one by one and does not support inquiry-based sensemaking of multiple taint flows and the impact of user choices and 3rd-party library models.
@@ 3.2. Soufflé and Logic Query Implementation
Once loaded into Soufflé, the questions in Table 1 are encoded as logic queries shown in Figures 2 and 3. For example, WhyFlow query uses transitive closure over edge to find whether data can reach a sink from a source, intersecting with library_flow to highlight which third party library’s APIs (orange nodes) appear along the path. WhyNotFlow similarly checks for “broken” transitive paths and identifies library_flow edges that act as sanitizers, effectively terminating any flow.
Supporting Divergent Sources/Sinks and Global Impact.
Queries like DivergentSinks or DivergentSources use logic rules to detect the last common node or first common node in two paths. We do so by computing the intersection of reachable sets for each source/sink pair. Meanwhile, GlobalImpact queries pivot on counting how frequently an API appears in distinct reachable flows; each API node is assigned a score via an aggregation rule over the factbase. The resulting integer is used to size the node (in the visualization) and thus show its “global impact.”
@@ 3.3. WhyFlow’s User Interface
WhyFlow supports multiple layout algorithms (e.g., breadth‐first, concentric). Users can reorganize the graph for improved clarity. A GlobalImpact query resizes API nodes by their number of occurrences on different flows.
@@ 4. User Study
Participants were asked to answer eight sensemaking questions about taint analysis results and how the configuration of third-party libraries impacts taint flows. Each user study task consists of eight questions in Table 3. The first six questions centered on why, why-not, and what-if questions about taint flows and the remaining two centered on quantifying taint flows.
@@
# of 3rd-party API functions 85
Table 2. Statistics of Taint Analysis Facts for Apache Dubbo.
@@ 4.1. Study Protocol
The study task is based on 383 taint warnings generated on Apache Dubbo with the CodeQL query shown in Listing 1 (codeql_externalapi_query, ). The resulting facts from this query are presented in Table 2.
@@ 4.1.2. Participants
We conducted a within-subject user study with a crossover design. We recruited 12 participants, including graduate students and professional developers from the industry. Their programming experience ranged from 1–3 years (4 participants), 4–6 years (3 participants), 7–10 years (3 participants), to over 10 years (2 participants). The average self-reported familiarity with taint analysis was 2 out of 5, suggesting that most participants had only moderate experience with this type of dataflow analysis.
7 participants are PhD students, 3 are MS students, 1 is an undergraduate and 1 is a professional developer from industry. We dropped one participant from the study, since they failed to complete half of both tasks. The average self-reported familiarity with taint analysis was 2.1 out of 5.
@@
Use of student participants.
While most of our study’s participants are students, prior work found that the findings based on student participants often generalize to a broader population as students have comparable performance to professionals in security-related tasks (sandoval_2023, ; acar_2016, ; acar_2017, ; salman_2015, ). Other studies (naiakshina_2017, ; naiakshina_2020, ; ko_2015, ) also emphasize that students can provide meaningful feedback when their expertise aligns with the tool’s target audience. Since WhyFlow is designed for end-user debugging, the use of student participants is methodologically sound. We also conducted a case study with two security professional participants, discussed in Section LABEL:casestudy. Also, the goal of WhyFlow is to help end-users who are not necessarily security experts. WhyFlow helps them to sensemake taint analysis results and debug them with our interrogative queries.
@@ 4.1.3. Protocol
Each participant took part in a 1.5-hour session. The study involved using both WhyFlow and the baseline CodeQL visualizer and the two tasks. The order of assigned tool (WhyFlow first vs. CodeQL first) and the assigned task (Problem Set #1 and Problem Set #2) was counterbalanced across participants through random assignment. We gave a 3-minute pre-study survey to collect background information, followed by a tutorial.
Next, each participant proceeded to the two tasks, each lasting 20 minutes. Each task had 8 questions that participants were tasked to answer. They had the option of skipping over questions and ending the task early. After each task, participants filled out post-task survey, which included the NASA TLX (5 minutes total). In the final 12-minute post-study survey, we asked participants to compare WhyFlow and CodeQL, and rate WhyFlow’s features.
@@ 4.1.4. Tutorial
We conducted a tutorial session to introduce participants to the functionality of both WhyFlow and the CodeQL VSCode plugin. The tutorials were structured around realistic dataflow scenarios. For WhyFlow, we guided users through all six query types. This involved selecting relevant sources, sinks, or API calls from dropdown menus and interpreting the graphical output rendered by WhyFlow. Participants were encouraged to hover over nodes to view fully qualified names, click on nodes to navigate directly to code locations, and adjust the layout using built-in options. Through these interactions, participants familiarized themselves with WhyFlow’s features.
We also conducted a tutorial session on using CodeQL’s visualizer. Participants learned how the plugin displays query results in a tabular list. We walked them through the process of expanding or collapsing the flows presented in the visualizer to inspect the intermediate steps between source to sink.
Finally, we explained the role of CodeQL “models” of external libraries, highlighting their role in the analysis. We reinforced core concepts: sources, sinks, sanitizers- and provided examples.
@@ 4.1.5. Tasks
Each task had a different set of questions (Problem Set #1 and Problem Set #2). The questions and their answers are shown in Table 3. To minimize learning effects, we balanced the order of tools that participants used first. Each task contained eight questions about ’why’, ’why-not’, ’what-if’, shown in Table 1 and the other two about quantitative aspects of taint flows (e.g., how many taint paths exist or how many third-party APIs appear in the taint paths). The questions were multiple-choice questions, some had multiple answers. We granted partial credit if participants selected some correct answers, but missed or added incorrect choices.
Program Point Set #1 Program Point Set #2
(1) Explain why a taint flow is permitted from getRequestURI(…) in [...].PageServlet.java to charAt(…) in [...].StringUtils.java. Name a third-party API permitting the flow. Answer: java.lang.String.substring(int beginIndex, int endIndex) (1) Explain why a taint flow is permitted from msg : String in [...].TelnetProcessHandler.java to json : String in [...].FastJsonImpl.java. Name a third-party API permitting the flow. Answer: java.lang.String.substring(...)
(2) Explain why no taint flow is permitted from msg : HttpRequest to warn(…) in [...].HttpProcessHandler.java. Answer: io.netty.handler.codec.http.HttpRequest.method(...) (2) Explain why no taint flow is permitted from msg : Http2StreamFrame in [...].TripleHttp2ClientResponseHandler.java to release(…) in [...].TripleClientStream.java. Answer: io.netty.handler.codec.http2. Http2DataFrame.isEndStream()
(3) Explain what sinks would no longer be reachable, if io.netty.handler.codec.http.HttpRequest.uri() were modeled as a sanitizer, starting from source msg : HttpRequest in [...].HttpProcessHandler.java. Answer: valueList in [...].HttpCommandDecoder.java and msg in [...].Log4jLogger.java (3) Explain what sinks would be no longer reachable, if io.netty.handler.codec.http2.Http2HeadersFrame.headers() is marked as a sanitizer, starting from source msg : Object in [...].TripleHttp2FrameServerHandler.java. Answer: path: [...].TriplePathResolver.java and headers: [...].TripleIsolationExecutorSupport.java
(4) Identify the program point that affects multiple taint flows ending at two sinks: sinks: path : String in [...].TriplePathResolver.java at line 41 and path : String in [...].TriplePathResolver.java at line 46, source: msg : Http2StreamFrame in [...].TripleHttp2ClientResponseHandler.java. Answer: toString(...) : String in [...].TripleServerStream.java (4) Identify the program point that affects multiple taint flows ending at two sinks: sinks: path : String in [...].TriplePathResolver.java – line 41 and path : String in [...].TriplePathResolver.java – line 46, source: msg : Object in [...].TripleHttp2FrameServerHandler.java. Answer: toString(...) : String in [...].TripleServerStream.java
(5) Identify the intermediary program point that influences multiple flows originating from input : ByteBuf in [...].NettyCodecAdapter.java and in : ByteBuf in [...].NettyPortUnificationServerHandler.java, ending at buffer : ByteBuf in [...].NettyBackedChannelBuffer.java. Answer: parameter this : NettyBackedChannelBuffer [buffer] : ByteBuf in [...].NettyBackedChannelBuffer.java (5) Identify the intermediary program point that influences multiple flows originating from msg : Http2StreamFrame in [...].TripleHttp2ClientResponseHandler.java and msg : Object in [...].TripleHttp2FrameServerHandler.java, ending at path : String in [...].TriplePathResolver.java. Answer: headers : Http2Headers in [...].TripleServerStream.java
(6) Which third-party APIs could have the most influence on the taint path from msg : Object in [...].TripleHttp2FrameServerHandler.java to path : String in [...].TriplePathResolver.java? Rank in the order of importance. Answer:
(1)io.netty.handler.codec.http2.Http2HeadersFrame.headers() (2)java.lang.CharSequence.toString()
@@
(7) Determine the number of pass-through API points from getRequestURI(…) in [...].PageServlet.java to charAt(…) in [...].StringUtils.java. Answer: 21 (7) Determine the number of pass-through API points from msg : Http2StreamFrame in [...].TripleHttp2ClientResponseHandler.java to path : String in [...].TriplePathResolver.java. Answer: 21
(8) Count how many different dataflow paths exist from in : ByteBuf in [...].NettyPortUnificationServerHandler.java to buffer : ByteBuf in [...].NettyBackedChannelBuffer.java. Answer: 2 (8) Count how many different dataflow paths exist from msg : Http2StreamFrame in [...].TripleHttp2ClientResponseHandler.java to path : String in [...].TriplePathResolver.java. Answer: 4
Table 3. Study Tasks: Program Sets #1 and #2 with their correct answers. Participants were assigned one of the two tasks to complete with CodeQL visualizer and the other with WhyFlow.
@@ 4.1.6. Criteria for Study Tasks
We designed questions based on realistic scenarios involving the analysis of flows from sources to sinks. These questions correspond to “why”, “why-not”, and “what-if” questions that developers may ask about a taint analysis or assess potential changes to models of external libraries. The tasks require pinpointing of sanitizers and pass-through APIs and assessing how many flows or nodes are present.
@@
After completing each task, participants completed a brief post-task questionnaire, which included the NASA TLX (hart1988development, ) (e.g., mental demand, time pressure, perceived success, effort, and frustration) and described their strategies to find answers to the questions. At the end of the session, participants rated user-interface features, confidence, and ease of use.

### v3

@@ 3. Speculative Analysis for Inquiry-based Debugging
WhyFlow enables inquiry-based debugging of spurious flows or missing flows. Figure 7 shows the precomputation of the taint facts by executing both the original CodeQL taint query and a less restricted version of this query (Section 3.1). After this, we convert the taint analysis results into Soufflé facts and store them in a “Results Database” (Section 3.2). Subsequently, WhyFlow queries this database to provide answers to the template questions (Section 3.3). We rely on several assumptions.
Assumption 1. We presume that a user of WhyFlow is familiar with the program they are inspecting, meaning that they can recognize that some flows are unexpected flows and some expected flows are missing, serving as a starting point for end-user debugging.
Assumption 2. We leverage CodeQL to obtain taint analysis results as is, and then we utilize Soufflé to run the interrogative debugging queries on top of CodeQL results.
Assumption 3. We assume that CodeQL results are created with the information flow models of third-party libraries, which is typical for modern taint analysis. We assume that inaccuracies in these plug-in models should be debugged, when a user suspects a missing flow or a spurious flow.
The goal of WhyFlow is not to disambiguate false positives from false negatives. Rather, WhyFlow helps with end-user debugging, explaining how configuration choices, such as source/sink definitions and third-party models, may lead to spurious or missing paths.
@@ 3.1. CodeQL
CodeQL is an open-source static analysis framework. Users can use QL, which is a declarative, object-oriented Domain-Specific Language (DSL) specifically designed for writing CodeQL queries to detect potential security vulnerabilities or other issues (codeql_docs, ).
A key factor in CodeQL’s comprehensiveness lies in its third-party library modeling. Each library API (including third-party dependencies) can be modeled to reflect how data flows through its methods. As manual creation of such models is time-consuming and error-prone, the CodeQL team has explored machine-learning-based techniques to infer library behaviors (nachtigall2022large, ). Nonetheless, incomplete or incorrect modeling remains a practical challenge.
Rather than relying on specific third-party library models and source and sink definitions, a general query may relax these configuration assumptions, revealing the “maximal” set of reachable flows from all possible sinks to all possible sources, ignoring third-party library models. Such an approach can help users uncover paths that would otherwise be missed due to incomplete or inaccurate library modeling. Moreover, comparing results from both model-aware and “maximal” queries can expose discrepancies, hints that the current model assumptions may need refinement. Our work builds on these insights by introducing an interactive, interrogative debugger, where users can easily toggle assumptions about third-party methods with pre-defined template queries.
Figure 8. CodeQL’s tree-based view shows each warning one by one and does not support inquiry-based sensemaking of multiple taint flows and the impact of user choices and third-party library models.
@@ 3.2. Soufflé and Logic Query Implementation
Once loaded into Soufflé, the questions in Table 1 are encoded as logic queries shown in Figures 2 and 3. For example, WhyFlow query uses transitive closure over edge to find whether data can reach a sink from a source, intersecting with library_flow to highlight which third-party library’s APIs (orange nodes) appear along the path. WhyNotFlow similarly checks for “broken” transitive paths and identifies library_flow edges that act as sanitizers, which terminate flows.
Supporting “DivergentSources/Sinks” and “GlobalImpact”.
DivergentSinks and DivergentSources queries use logic rules to detect the last common node or first common node in two paths. They compute the intersection of reachable sets for each source/sink pair. Meanwhile, GlobalImpact queries count how frequently an API appears in distinct reachable flows; each API node is assigned a score via an aggregation rule. WhyFlow sizes the node (in the visualization) based on its score, visualizing its “global impact.”
@@ 3.3. WhyFlow’s User Interface
WhyFlow supports multiple layout algorithms (e.g., breadth‐first, concentric). A GlobalImpact query resizes API nodes by their number of occurrences on different flows.
@@ 4. User Study
Participants were asked to answer eight sensemaking questions about taint analysis results and how the configuration of third-party libraries impacts taint flows. Each user study task consists of eight questions shown in Table 3. The first six questions are centered on why, why-not, and what-if questions about taint flows and the remaining two are focused on quantifying taint flows.
@@
# of third-party API functions 85
Table 2. Statistics of Taint Analysis Facts for Apache Dubbo (Kevin2025apache, ).
@@ 4.1. Study Protocol
The study task is based on 383 taint warnings generated on Apache Dubbo (Kevin2025apache, ) with the CodeQL query shown in Listing 1 (codeql_externalapi_query, ). The resulting facts from this query are presented in Table 2.
@@ 4.1.2. Participants
We conducted a within-subject user study with a crossover design. We recruited 12 participants, including graduate students and professional developers from the industry. Their programming experience ranged from 1–3 years (4 participants), 4–6 years (3 participants), 7–10 years (3 participants), to over 10 years (2 participants). The average self-reported familiarity with taint analysis was 2.1 out of 5, suggesting that most participants had only moderate experience with this type of dataflow analysis.
7 participants are PhD students, 3 are MS students, 1 is an undergraduate and 1 is a professional developer from industry. We dropped one participant from the study, since the participant failed to complete half of both tasks.
@@ 4.1.3. Protocol
Each participant took part in a 1.5-hour session. The study involved using both WhyFlow and the baseline CodeQL visualizer for the two tasks. The order of assigned tool (WhyFlow first vs. CodeQL first) and the assigned task (Problem Set #1 and Problem Set #2) was counterbalanced across participants through random assignment. We gave a 3-minute pre-study survey to collect background information, followed by a tutorial.
Next, each participant proceeded with the two tasks, each lasting 20 minutes. Each task had 8 questions that participants were tasked to answer. They had the option of skipping over questions and ending the task early. After each task, participants filled out a 5-minute post-task survey, which included the NASA-TLX questions (hart1988development, ) (e.g., mental demand, time pressure, perceived success, effort, and frustration). In the final 12-minute post-study survey, we asked participants to compare WhyFlow and CodeQL, and rate WhyFlow’s features.
@@ 4.1.4. Tutorial
We conducted a tutorial session to introduce participants to the functionality of both WhyFlow and the CodeQL VSCode plugin. For WhyFlow, we guided users through all six query types. This involved selecting relevant sources, sinks, or API calls from dropdown menus and interpreting the graphical output rendered by WhyFlow. Participants were encouraged to interact with WhyFlow’s features to familiarize themselves with them.
We also conducted a tutorial session on using CodeQL’s visualizer. We walked the participants through the process of viewing query results in a tabular list, and expanding or collapsing the flows to inspect the intermediate steps between source and sink.
Finally, we explained the role of CodeQL “models” of external libraries, highlighting their role in the analysis. We reinforced core concepts—sources, sinks, and sanitizers—and provided examples.
@@ 4.1.5. Tasks
Each task had a different set of questions (Problem Set #1 and Problem Set #2). The questions and their answers are shown in Table 3. Each task contained eight questions about ‘why’, ‘why-not’, ‘what-if,’ as shown in Table 1, and two additional questions about the quantitative aspects of taint flows (e.g., how many taint paths exist or how many third-party APIs appear in the taint paths). The questions were multiple-choice questions, some having multiple answers. We granted partial credit if participants selected some correct answers, but missed or added incorrect choices.
Problem Set #1 Problem Set #2 (1) Explain why a taint flow is permitted from getRequestURI(...) in [...].PageServlet.java to charAt(…) in [...].StringUtils.java. Name a third-party API permitting the flow. Answer: java.lang.String.substring(int beginIndex, int endIndex) (1) Explain why a taint flow is permitted from msg: String in [...].TelnetProcessHandler.java to json: String in [...].FastJsonImpl.java. Name a third-party API permitting the flow. Answer: java.lang.String.substring(...) (2) Explain why no taint flow is permitted from msg: HttpRequest to warn(...) in [...].HttpProcessHandler.java. Answer: io.netty.handler.codec.http.HttpRequest.method(...) (2) Explain why no taint flow is permitted from msg: Http2StreamFrame in [...].TripleHttp2ClientResponseHandler.java to release(…) in [...].TripleClientStream.java. Answer: io.netty.handler.codec.http2. Http2DataFrame.isEndStream() (3) Explain what sinks would no longer be reachable, if io.netty.handler.codec.http.HttpRequest.uri() were modeled as a sanitizer, starting from source msg: HttpRequest in [...].HttpProcessHandler.java.[1ex] Answer: valueList in [...].HttpCommandDecoder.java and msg in [...].Log4jLogger.java (3) Explain what sinks would be no longer reachable, if io.netty.handler.codec.http2.Http2HeadersFrame.headers() is marked as a sanitizer, starting from source msg: Object in [...].TripleHttp2FrameServerHandler.java. Answer: path: [...].TriplePathResolver.java and headers: [...].TripleIsolationExecutorSupport.java (4) Identify the program point that affects multiple taint flows ending at two sinks: sinks: path: String in [...].TriplePathResolver.java at line 41 and path: String in [...].TriplePathResolver.java at line 46, source: msg: Http2StreamFrame in [...].TripleHttp2ClientResponseHandler.java. Answer: toString(...): String in [...].TripleServerStream.java (4) Identify the program point that affects multiple taint flows ending at two sinks: sinks: path: String in [...].TriplePathResolver.java – line 41 and path: String in [...].TriplePathResolver.java – line 46, source: msg: Object in [...].TripleHttp2FrameServerHandler.java. Answer: toString(...): String in [...].TripleServerStream.java (5) Identify the intermediary program point that influences multiple flows originating from input: ByteBuf in [...].NettyCodecAdapter.java and in: ByteBuf in [...].NettyPortUnificationServerHandler.java, ending at buffer: ByteBuf in [...].NettyBackedChannelBuffer.java. Answer: parameter this : NettyBackedChannelBuffer [buffer]: ByteBuf in [...].NettyBackedChannelBuffer.java (5) Identify the intermediary program point that influences multiple flows originating from msg: Http2StreamFrame in [...].TripleHttp2ClientResponseHandler.java and msg: Object in [...].TripleHttp2FrameServerHandler.java, ending at path: String in [...].TriplePathResolver.java. Answer: headers: Http2Headers in [...].TripleServerStream.java (6) Which third-party APIs could have the most influence on the taint path from msg: Object in [...].TripleHttp2FrameServerHandler.java to path: String in [...].TriplePathResolver.java? Rank in the order of importance. Answer:
(1) io.netty.handler.codec.http2.Http2HeadersFrame.headers()
(2) java.lang.CharSequence.toString()
@@
(7) Determine the number of pass-through API points from getRequestURI(...) in [...].PageServlet.java to charAt(...) in [...].StringUtils.java. Answer: 21 (7) Determine the number of pass-through API points from msg: Http2StreamFrame in [...].TripleHttp2ClientResponseHandler.java to path: String in [...].TriplePathResolver.java. Answer: 21 (8) Count how many different dataflow paths exist from in: ByteBuf in [...].NettyPortUnificationServerHandler.java to buffer: ByteBuf in [...].NettyBackedChannelBuffer.java. Answer: 2 (8) Count how many different dataflow paths exist from msg: Http2StreamFrame in [...].TripleHttp2ClientResponseHandler.java to path: String in [...].TriplePathResolver.java. Answer: 4
Table 3. Study Tasks: Problem Sets #1 and #2 with their correct answers. Participants were assigned one of the two tasks to complete with CodeQL visualizer and the other with WhyFlow.
@@ 4.1.6. Criteria for Study Tasks
We designed questions based on realistic scenarios involving the analysis of taint flows. These questions correspond to “why”, “why-not”, and “what-if” questions that developers may ask about a taint analysis or assess potential changes to models of the external libraries.
@@
After completing each task, participants completed a brief post-task questionnaire, which included the NASA-TLX questions. They also described their strategies to find answers to the questions in this questionnaire. At the end of the session, participants rated user-interface features, confidence, and ease of use.